OpenStudy (anonymous):

For a certain experiment, we were able to obtain five samples with mean = 25. Assume that the sample is coming from a normal distribution with o = 9(o is the greek symbol, forget the name. Compute 95% upper confidence bound for the population mean mew. I keep getting 32.889. Want to check to see if thats right.

OpenStudy (kropot72):

The sample mean x bar = 25 The sample size n = 5 The population standard deviation sigma = 9 The 95% upper confidence bound for the population mean is found from: \[xbar+1.96\frac{\sigma}{\sqrt{n}}=25+(1.96\frac{9}{\sqrt{5}})\] This calculates out to the same result as you get.